A reaction occurs where 7.0x 10 12 atoms of hydrogen gas (H2) combine with 3.5 x10 12 atoms of graphite. (C) Which molecular model is inconsistent with these data?  

C3H12 CH2 CH4 C2H8 (My math says the answer should have a 1:3 ratio which means none of these are right??)
Answer #1

Number of H- atoms = 7 x 1012

Number of moles of H- atoms = (7 x 1012)/(6.022 x 1023) = 1.16 x 10-11 mol

Number of C-atoms = 3.5 x 1012

Number of moles of C- atoms = (3.5 x 1012)/(6.022 x 1023) = 5.81 x 10-12 mol

Moles ratios of H and C = 1.16 x 10-11 : 5.81 x 10-12 = 2 : 1

Therefore, there are 2 moles of H-atoms and 1 mol of Carbon atoms. Hence, the molecular model consistent with these data is CH2.